# RL Zero: Zero-Shot Language to Behaviors without any supervision

### v1

@@ 5.3 Ablation and Failure Cases
1. What I cannot imagine, I cannot imitate: The video generation model used in our work from Mazzaglia et al. (2024) is fairly small and limited in capability. We encountered limitations when generating complex behaviors with this model and found it to be sensitive to prompt engineering. Fortunately, as models get bigger and are trained on a larger set of data, this limitation can be overcome. Figure 6a shows some examples of these failures with the corresponding prompts.
@@ 6 Conclusion
Language presents an appealing and human-friendly alternative to reward design for task specification. In this work, we presented a completely unsupervised approach for grounding language to low-level behavior in a zero-shot manner. A completely unsupervised approach allows us to bypass requiring costly annotators for labeling a wide variety of behaviors with language, and a zero-shot approach allows us to avoid training during deployment time along with the advantage of generating the behaviors instantaneously. We propose RLZero, a framework to imagine what a behavior specified by a text prompt looks like and to ground that imagination to a policy via imitation. Unlike reward functions, this approach is not prone to reward hacking as the distribution matching objective specifies the task completely and accurately. Our evaluations show that the behaviors generated by RLZero show an improvement over using reward functions derived from image-language of video-language models.
Future Directions: RLZero opens up the possibility of prompting to generate a policy. Zero-shot approaches are always expected to be near-optimal due to the projection of a reward to a low dimensional space as well as limited coverage of offline interaction data. But this serves as a good initialization for further fine-tuning. How to fine-tune efficiently without forgetting remains an open question. Furthermore, learned skills can be combined according to the hierarchy specified in language instructions, allowing for the completion of complex long-horizon tasks. Since the mechanism of RLZero allows for interoperability to some extent by observing the nearest states as well as imagination, automatic failure detection becomes appealing. For the setting of prompt-to-policy, we lack accurate evaluation metrics since the true reward function is unknown, and human evaluation can be subjective. Finally, with larger context-window video understanding models, we believe an end-to-end pipeline of language embedding to task embedding (imagination-free RLZero) can become more appealing.
@@ Appendix A Proof For Theorem 1
The first term is an RL objective with a reward function given by $\log(\frac{\rho^{E}}{\rho})$, and the second term is an offline regularization to constrain the behaviors of offline datasets. Following prior works Kim et al. (2022); Ma et al. (2022a), since our BFM is trained on an offline dataset and limited to output skills in support of dataset actions, and we can ignore the regularization to infer the latent $z$ parameterizing the skill. A heuristic yet performant alternative is to use a shaped reward function of $\frac{\rho^{E}}{\rho}$, which allows us to avoid training the discriminator completely and was shown to lead to performant imitation in Pirotta et al. (2023). ∎
@@ B.1.1 DM-control environments
We use continuous control environments from the DeepMind Control Suite (Tassa et al., 2018).
Walker: It has a 24 dimensional state space consisting of joint positions and velocities and 6 dimensional action space where each dimension of action lies in $[-1,1]$. The system represents a planar walker.
Cheetah: It has a 17 dimensional state space consisting of joint positions and velocities and 6 dimensional action space where each dimension of action lies in $[-1,1]$. The system represents a planar biped “cheetah”.
Quadruped: It has a 78 dimensional state space consisting of joint positions and velocities and 12 dimensional action space where each dimension of action lies in $[-1,1]$. The system represents a 3-dimensional ant with 4 legs.
Stickman: Stickman was recently introduced as a task that bears resemblance to a humanoid in Mazzaglia et al. (2024). It has a 44 dimensional observation space and a 10 dimensional action space where each dimension of action lies in $[-1,1]$.
@@ B.2 Evaluation Protocol
For the set of task prompts specified by language, there is no ground truth reward function and there does not exist a reliable quantitative metric to verify which of the methods perform better. Instead, since humans communicate their intents via language, humans are the best judge of whether the agent has demonstrated the behavior they intended to convey. In this work we use a Multimodal LLM as a judge, following studies by prior works demonstrating the correlation of LLMs judgment to humans (Chen et al., 2024). We use GPT-4o model as the judge, where the GPT-4o model is provided with two videos, one generated by a base method, and another generated by one of the methods we consider, and asked for preference between which video is better explained by the text prompt for the task. When inputting the videos to the judge, we randomize the order of the baseline and proposed methods to reduce the effect of anchoring bias. The prompt we use to compare the two methods is given here:
@@ B.3 Dataset Collection for Zero-Shot RL
For Cheetah, Walker, Quadruped, and Stickman environments, our data is collected following a pure exploration algorithm with no extrinsic rewards. In this work, we use intrinsic rewards obtained from Random Network Distillation (Burda et al., 2018) to collect our dataset based on the protocol by ExoRL (Yarats et al., 2022) and using the implementation from repository ExoRL repository. For Cheetah, Walker, and Quadruped, our dataset comprises 5000 episodes and equivalently 5 million transitions, and for Stickman, our dataset comprises 10000 episodes or equivalently 10 million transitions. Due to the high dimensionality of action space in Stickman, RND does not discover a lot of meaningful behaviors; hence we additionally augment the dataset with 1000 episodes from the replay buffer of training for a ‘running’ reward function and 1000 episodes of replay buffer trained on a ‘standing’ reward function.
@@ B.4 Baselines
We implement two offline RL baselines to compare with RLZero– Implicit Q-learning (IQL, Kostrikov et al. (2021)) and Offline TD3 (TD3, Fujimoto & Gu (2021)). Both of these methods share the same offline dataset as used to learn the successor measure in RLZero, which is described in Section 5, and gathered using RND. Since these datasets are reward-free, we must still construct a reward function that provides meaningful rewards for an agent achieving the behavior that aligns with the text prompt. Formally, given language instruction $e^{l}\in\mathcal{E}^{l}$, frame stack $(o_{t-k},o_{t-k+1},\ldots,o_{t})\in\mathcal{I}$, and embedding VLM $\phi:\mathcal{E}\rightarrow\mathcal{Z}$, which can also embed frame stacks $\phi:\mathcal{I}\rightarrow\mathcal{Z}$ (and where observations $o_{i}\in\mathcal{I}$, and we use $o_{i}$ for this section), the reward for a corresponding language instruction and frame stack $k$ is the cosine similarity between the stacked language embedding and the frame embedding:
@@
Image-language reward (SigLIP Zhai et al. (2023)): take a stack of 3 frames encode them using SigLIP, then the reward is computed as the cosine distance of the embeddings and the SigLIP embedding of language.
Video-language reward (InternVideo2 Wang et al. (2024a)): this method takes in previous frames $o_{0:t-1}$ as context and uses it to generate an embedding of the current frame observation $o_{t}$. The video encoder then takes the cosine similarity of $\phi(o_{0:t})$ and $\phi(e^{l})$. This allows the reward function to provide rewards based not only on reaching certain states, but the agent exhibiting temporally extended behaviors that match the behavior. In practice, providing rewards using an image-based encoder for frame stacks can be challenging for tasks such as walking because they require context, and video-based rewards offer a way to better encode the temporal context.
@@ B.4.1 Offline RL
Implicit Q-learning (Kostrikov et al., 2021) Implicit Q-learning builds on the classic TD error (revised in our context of language-instruction rewards):
@@
TD3 (Fujimoto et al., 2018):
TD3 was demonstrated to be the best performing algorithm when learning from exploratory RND datasets in (Yarats et al., 2022). While TD3 does not explicitly address the challenges discussed in implicit Q-learning and learns using Bellman Optimality backups, the approach is simple and works well in practice. The algorithm uses a deterministic policy extraction $\pi:\mathcal{S}\rightarrow\mathcal{A}$ to give the following objective:
@@ B.5.1 Text to Imagined Behavior with Video models
To generate a proposed video frame sequence, we utilize the GenRL architecture and provide the workflow using equations from the original paper (Mazzaglia et al., 2024). First, the desired text prompt is embedded with the underlying video foundation model InternVideo2 (Wang et al., 2024a) $e^{(l)}=f_{PT}^{(l)}(y)$. These embeddings are then repeated $n_{frames}$ times (we use $n_{frames}=32$) to match the temporal structure expected by the world model. The repeated text embeddings are passed through an aligner module $e(v)=f_{\psi}(e^{(l)})$. The aligner is implemented as a UNet and it is used to address the multimodality gap (Liang et al., 2022) when embeddings from different modalities occupy distinct regions in the latent space. Next, actions are constructed by concatenating the aligned video embeddings with temporal embeddings. The temporal embeddings are one-hot encodings of the time step modulo $n_{frames}$ providing frame-level positional information. The first action is passed to the world model connector $p_{\psi}(s_{t}|e)$ to initialize the latent state. For each subsequent time step, the sequence model $h_{t}=f_{\phi}(s_{t-1},a_{t-1},h_{t-1})$ (implemented as a GRU) updates the deterministic state $h_{t}$. The deterministic state $h_{t}$ is mapped to a stochastic latent state ($s_{t}$) using the dynamics predictor $p_{\phi}(s_{t}|h_{t})$. The dynamics predictor, implemented as an ensemble of MLPs, predicts the sufficient statistics (mean and standard deviation) for a Normal distribution over $s_{t}$. During inference, the mean of this distribution is used as the latent state. Finally, the latent state $s_{t}$ is passed to the convolutional decoder $p_{\phi}(x_{t}|s_{t})$ to reconstruct the video frame $x_{t}$. This process is repeated for all time steps $(t=1,...,n_{frames})$.
@@ B.5.2 Grounding imagined observations to observations in offline dataset
As described in Section 4, we ground imagined sequences by retrieving real offline states based on similarity in an embedding space. This enables us to create a suitable $z$-vector for distribution matching which is the expected value of the state features under the distribution of imagined states ($\rho_{imagined})$. During our dataset collection phase, we save both the agent’s proprioceptive state as well as the corresponding rendered images and search over the images to then find the corresponding state. Our code supports both stacked-frame embeddings and single-frame embeddings. We find that stacked-frame embeddings were helpful in modeling temporal dependencies through velocity and acceleration, which are crucial for recreating the intended behavior. SigCLIP (Zhai et al., 2023), which replaces CLIP’s (Radford et al., 2021a) softmax-based contrastive loss with a pairwise sigmoid loss, resulted in qualitatively better matches to exact positions within sequences, imitating behavior more accurately than CLIP. For both models, we use the OpenCLIP (Ilharco et al., 2021) framework. Our matching process first involves precomputing embeddings offline, which are stored in chunks of up to 100,000 frames to optimize memory usage and retrieval speed. During inference, we load this file and embed the query frame sequence from GenRL (Mazzaglia et al., 2024) into the same latent space. We process these query embeddings by dividing them into chunks of $k$-frame sequences ($k$ generally 3 or 5), where each sequence consists of the current frame and the $k-1$ preceding frames. If there are not enough preceding frames, we repeat the first frame to fill the gap. For each chunk of saved embeddings, we compute dot products between the query chunk and all subsequences of size $k$ in the saved embeddings. We track the highest similarity score for each query chunk and return the frames corresponding to the closest embedding sequences.
@@ B.5.3 Training a zero-shot RL agent
In this work, we chose Forward-Backward (FB) (Touati & Ollivier, 2021) as our zero-shot RL algorithm and train it on proprioceptive inputs. Our implementation follows closely from the authors codebase . Specifically, FB trains Forward, Backward, and Actor networks. The backward networks are used to map a demonstration or a reward function to a skill, which is then used to learn a latent-conditional Actor. The hyperparameters for our FB implementation are listed below:
Implementation: We build upon the codebase for FB https://github.com/facebookresearch/controllable_agent and implement all the algorithms under a uniform setup for network architectures and same hyperparameters for shared modules across the algorithms. We keep the same method agnostic hyperparameters and use the author-suggested method-specfic hyperparameters. The hyperparameters for all methods can be found in Table 2:
@@ B.6 Imagination-Free RLZero
Pretraining: We first generate a dataset of episodes containing diverse behaviors by rolling out the behavior foundation model conditioned on a uniformly random sampled $z$. The resulting image observation sequences are then down-sampled (by 8) and sliced to break up each episode into smaller chunks of length 8; this preprocessing step helps increase the behavioral diversity and improves the ability of the ViFM to capture semantic meaning. The resulting clips are then embedded using a ViFM (InternVideo2 (Wang et al., 2024a)) where each embedding is denoted by $e$ (as in Section 5.3). Now we have a set of sequences of length 8 consisting of image observation along with their proprioceptive states, and the embedding for the image sequence.
@@
Inference: During inference, the language prompt is embeddedto a latent vector $e^{l}$. A known issue with multimodal embedding models is the embedding gap (Liang et al., 2022), which makes the video embeddings unaligned with text embeddings. To account for this gap, we use an aligner trained in an unsupervised fashion from previous work Mazzaglia et al. (2024) to align the language embedding ($e^{l}_{aligned}$). Then the aligned embedding is passed through the latent mapper to get the policy conditioning $z_{imitation}$ which gives us the policy that achieves the desired behavior specified through language.
@@ Appendix C Additional Results
C.1 Zero-shot imitation: Discriminator vs Discriminator-free

### v2

@@ 5.3 Ablation and Limitations
Imagination-free behavior generation: While the imagine, project, and imitate framework allows for the interpretability of the agent’s behavior, we investigate if we can amortize the imagination and embedding search cost by directly mapping the language embedding to the skill embedding in the BFM’s latent space. For this, we consider sampling $z$ uniformly in the latent space of the BFM and embedding the generated image observation sequence through a video-language encoder, which we denote by $e$. Given the observation sequence, we generate the $z_{imit}$ using the zero-shot inference process and learn a mapping from $e\rightarrow z_{imit}$ using a small 3-layer MLP. On the same tasks considered in Fig 3, we observe imagination-free RLZero to have a win rate of 65.71% over TD3 base model on Walker environment when compared to RLZero that had a win rate of $91.4\%$ (detailed results in Table 7). A thorough explanation of imagination-free RLZero can be found in Appendix B.6.
1. What I cannot imagine, I cannot imitate: The video generation model used in our work is fairly small and limited in capability. We encountered limitations when generating complex behaviors with this model and found it to be sensitive to prompt engineering. Fortunately, as models get bigger and are trained on a larger set of data, this limitation can be overcome. Figure 4a shows some examples of these failures with the corresponding prompts.
@@ 6 Conclusion
Specifying desired behavior to a learned agent is a longstanding problem in sequential decision-making. In this work, we take a step towards leveraging intuitive mediums such as language and cross-embodiment to specify desired policies. The RLZero framework bypasses costly annotations, reward hacking, and deployment time training by leveraging the rich capabilities of behavioral foundation models. Our introduced framework demonstrates a technique for directly translating from intuitive mediums for task specification to the latent space of policies without any test-time training. We demonstrate empirically that RLZero not only recovers policies with zero gradient steps at evaluation time, but these policies also achieve qualitative and quantitative improvement over prior baselines that leverage VLMs as a source or reward or perform test-time training.
@@ Appendix A Proof For Theorem 1
The first term is an RL objective with a reward function given by $\log(\frac{\rho^{E}}{\rho})$, and the second term is an offline regularization to constrain the behaviors of offline datasets. Following prior works [35, 46], since our BFM is trained on an offline dataset and limited to output skills in support of dataset actions, and we can ignore the regularization to infer the latent $z$ parameterizing the skill. A heuristic yet performant alternative is to use a shaped reward function of $\frac{\rho^{E}}{\rho}$, which allows us to avoid training the discriminator completely and was shown to lead to performant imitation in [57]. ∎
@@ B.1.1 DM-control environments
We use continuous control environments from the DeepMind Control Suite [72].
Walker: The agent has a 24 dimensional state space consisting of joint positions and velocities and 6 dimensional action space where each dimension of action lies in $[-1,1]$. The system represents a planar walker.
Cheetah: The agent has a 17 dimensional state space consisting of joint positions and velocities and 6 dimensional action space where each dimension of action lies in $[-1,1]$. The system represents a planar biped “cheetah".
Quadruped: The agent has a 78 dimensional state space consisting of joint positions and velocities and 12 dimensional action space where each dimension of action lies in $[-1,1]$. The system represents a 3-dimensional ant with 4 legs.
Stickman: Stickman was recently introduced as a task that bears resemblance to a humanoid in [49]. It has a 44 dimensional observation space and a 10 dimensional action space where each dimension of action lies in $[-1,1]$.
SMPL 3D Humanoid: The agent has a 358 dimensional state space consisting of joint positions and velocities and 69 dimensional action space where each dimension of action lies in $[-1,1]$. The system represents a 3-dimensional humanoid.
@@ B.2 Evaluation Protocol
For the set of task prompts specified by language, there is no ground truth reward function and there does not exist a reliable quantitative metric to verify which of the methods perform better. Instead, since humans communicate their intents via language, humans are the best judge of whether the agent has demonstrated the behavior they intended to convey. In this work we use a Multimodal LLM as a judge, following studies by prior works demonstrating the correlation of LLMs judgment to humans [14]. We use GPT-4o model as the judge, where the GPT-4o model is provided with two videos, one generated by a base method, and another generated by one of the methods we consider, and asked for preference between which video is better explained by the text prompt for the task. When inputting the videos to the judge, we randomize the order of the baseline and proposed methods to reduce the effect of anchoring bias. The prompt we use to compare the two methods is given here:
@@ B.3 Dataset Collection for Zero-Shot RL
For Cheetah, Walker, Quadruped, and Stickman environments, our data is collected following a pure exploration algorithm with no extrinsic rewards. In this work, we use intrinsic rewards obtained from Random Network Distillation [10] to collect our dataset based on the protocol by ExoRL [86] and using the implementation from repository ExoRL repository. For Cheetah, Walker, and Quadruped, our dataset comprises 5000 episodes and equivalently 5 million transitions, and for Stickman, our dataset comprises 10000 episodes or equivalently 10 million transitions. Due to the high dimensionality of action space in Stickman, RND does not discover a lot of meaningful behaviors; hence we additionally augment the dataset with 1000 episodes from the replay buffer of training for a ‘running’ reward function and 1000 episodes of replay buffer trained on a ‘standing’ reward function.
@@ B.4 Baselines
We implement two offline RL baselines to compare with RLZero– Implicit Q-learning (IQL, [37]) and Offline TD3 (TD3, [22]). Both of these methods share the same offline dataset as used to learn the successor measure in RLZero, which is described in Section 5, and gathered using RND. Since these datasets are reward-free, we must still construct a reward function that provides meaningful rewards for an agent achieving the behavior that aligns with the text prompt. Formally, given language instruction $e^{l}\in\mathcal{E}^{l}$, frame stack $(o_{t-k},o_{t-k+1},\ldots,o_{t})\in\mathcal{I}$, and embedding VLM $\phi:\mathcal{E}\rightarrow\mathcal{Z}$, which can also embed frame stacks $\phi:\mathcal{I}\rightarrow\mathcal{Z}$ (and where observations $o_{i}\in\mathcal{I}$, and we use $o_{i}$ for this section), the reward for a corresponding language instruction and frame stack $k$ is the cosine similarity between the stacked language embedding and the frame embedding:
@@
Image-language reward (SigLIP [89]): take a stack of 3 frames encode them using SigLIP, then the reward is computed as the cosine distance of the embeddings and the SigLIP embedding of language.
Video-language reward (InternVideo2 [79]): this method takes in previous frames $o_{0:t-1}$ as context and uses it to generate an embedding of the current frame observation $o_{t}$. The video encoder then takes the cosine similarity of $\phi(o_{0:t})$ and $\phi(e^{l})$. This allows the reward function to provide rewards based not only on reaching certain states, but the agent exhibiting temporally extended behaviors that match the behavior. In practice, providing rewards using an image-based encoder for frame stacks can be challenging for tasks such as walking because they require context, and video-based rewards offer a way to better encode the temporal context.
@@ B.4.1 Offline RL
Implicit Q-learning [37] Implicit Q-learning builds on the classic TD error (revised in our context of language-instruction rewards):
@@
TD3 [23]:
TD3 was demonstrated to be the best performing algorithm when learning from exploratory RND datasets in [86]. While TD3 does not explicitly address the challenges discussed in implicit Q-learning and learns using Bellman Optimality backups, the approach is simple and works well in practice. The algorithm uses a deterministic policy extraction $\pi:\mathcal{S}\rightarrow\mathcal{A}$ to give the following objective:
@@ B.5.1 Text to Imagined Behavior with Video models
To generate a proposed video frame sequence, we utilize the GenRL architecture and provide the workflow using equations from the original paper [49]. First, the desired text prompt is embedded with the underlying video foundation model InternVideo2 [79] $e^{(l)}=f_{PT}^{(l)}(y)$. These embeddings are then repeated $n_{frames}$ times (we use $n_{frames}=32$) to match the temporal structure expected by the world model. The repeated text embeddings are passed through an aligner module $e(v)=f_{\psi}(e^{(l)})$. The aligner is implemented as a UNet and it is used to address the multimodality gap [40] when embeddings from different modalities occupy distinct regions in the latent space. Next, the aligned video embeddings are concatenated with temporal embeddings. The temporal embeddings are one-hot encodings of the time step modulo $n_{frames}$ providing frame-level positional information. The first embedding is passed to the world model connector $p_{\psi}(s_{t}|e)$ to initialize the latent state. For each subsequent time step, the sequence model $h_{t}=f_{\phi}(s_{t-1},a_{t-1},h_{t-1})$ (implemented as a GRU) updates the deterministic state $h_{t}$. The deterministic state $h_{t}$ is mapped to a stochastic latent state ($s_{t}$) using the dynamics predictor $p_{\phi}(s_{t}|h_{t})$. The dynamics predictor, implemented as an ensemble of MLPs, predicts the sufficient statistics (mean and standard deviation) for a Normal distribution over $s_{t}$. During inference, the mean of this distribution is used as the latent state. Finally, the latent state $s_{t}$ is passed to a convolutional decoder $p_{\phi}(x_{t}|s_{t})$ to reconstruct the video frame $x_{t}$. This process is repeated for all time steps $(t=1,...,n_{frames})$.
@@ B.5.2 Grounding imagined observations to observations in offline dataset
As described in Section 4, we ground imagined sequences by retrieving real offline states based on similarity in an embedding space. This enables us to create a suitable $z$-vector for distribution matching which is the expected value of the state features under the distribution of imagined states ($\rho_{imagined})$. During our dataset collection phase, we save both the agent’s proprioceptive state as well as the corresponding rendered images and search over the images to then find the corresponding state. Our code supports both stacked-frame embeddings and single-frame embeddings. We find that stacked-frame embeddings were helpful in modeling temporal dependencies through velocity and acceleration, which are crucial for recreating the intended behavior. SigLIP [89], which replaces CLIP’s [60] softmax-based contrastive loss with a pairwise sigmoid loss, resulted in qualitatively better matches to exact positions within sequences, imitating behavior more accurately than CLIP. For both models, we use the OpenCLIP [30] framework. Our matching process first involves precomputing embeddings offline, which are stored in chunks of up to 100,000 frames to optimize memory usage and retrieval speed. During inference, we load this file and embed the query frame sequence from GenRL [49] into the same latent space. We process these query embeddings by dividing them into chunks of $k$-frame sequences ($k$ is generally 3 or 5), where each sequence consists of the current frame and the $k-1$ preceding frames. If there are not enough preceding frames, we repeat the first frame to fill the gap. For each chunk of saved embeddings, we compute dot products between the query chunk and all subsequences of size $k$ in the saved embeddings. We track the highest similarity score for each query chunk and return the frames corresponding to the closest embedding sequences.
@@ B.5.3 Training a zero-shot RL agent
In this work, we chose Forward-Backward (FB) [76] as our zero-shot RL algorithm and trained it on proprioceptive inputs. Our implementation follows closely from the author’s codebase . Specifically, FB trains Forward, Backward, and Actor networks. The backward networks are used to map a demonstration or a reward function to a skill, which is then used to learn a latent-conditional Actor. Our experiments were performed on NVIDIA-A40 and AMD EPYC 7763 64-Core Processor machine. The hyperparameters for our FB implementation are listed below:
Implementation: We build upon the codebase for FB https://github.com/facebookresearch/controllable_agent and implement all the algorithms under a uniform setup for network architectures and the same hyperparameters for shared modules across the algorithms. We keep the same method-agnostic hyperparameters and use the author-suggested method-specific hyperparameters. The hyperparameters for all methods can be found in Table 3:
@@ B.6 Imagination-Free RLZero
Pretraining: We first generate a dataset of episodes containing diverse behaviors by rolling out the behavior foundation model conditioned on a uniformly random sampled $z$. The resulting image observation sequences are then down-sampled (by 8) and sliced to break up each episode into smaller chunks of length 8; this preprocessing step helps increase the behavioral diversity and improves the ability of the ViFM to capture semantic meaning. The resulting clips are then embedded using a ViFM (InternVideo2 [79]) where each embedding is denoted by $e$ (as in Section 5.3). Now, we have a set of sequences of length 8 consisting of image observation along with their proprioceptive states, and the embedding for the image sequence.
@@
Inference: During inference, the language prompt is embedded to a latent vector $e^{l}$. A known issue with multimodal embedding models is the embedding gap [40], which makes the video embeddings unaligned with text embeddings. To account for this gap, we use an aligner trained in an unsupervised fashion from previous work [49] to align the language embedding ($e^{l}_{aligned}$). Then the aligned embedding is passed through the latent mapper to get the policy conditioning $z_{imitation}$ which gives us the policy that achieves the desired behavior specified through language.
@@ Appendix C Additional Results
C.1 Visualizations
C.2 Zero-shot imitation: Discriminator vs Discriminator-free